A rectangle has length 3cm greater than its width, if it has an area of 28cm2, what is the dimesions of the rectangle??
Answer by ewatrrr(24785)

 
Hi,

Let x represent the width
(x+3cm)x = 28cm^2
x^2 + 3x -28 = 0 (Tossing out the negative solution for unit measure)
(x+7)(x-4) = 0 , x = 4cm, the width. The length is 7cm